ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: Semenov, E. M., and F. L. Hernández Disjoint Systems and Complemented Subspaces in Rearrangement Invariant Spaces The class N of all those rearrangement invariant spaces is studied on [0, ∞) for which all sublattices generated by disjoint shifts of a single element are complemented. The class N is stable with respect to the complex interpolation method but is not with respect to the real method. A Lorentz space L_{p,q} belongs to N iff 1 ≤ q ≤ p. Properties of the class N are investigated.